Fermi National Laboratory


Accelerator Update for May 2, 2005 - May 4, 2005

Monday May 2
The day shift began with Operations monitoring store 4122 and a small Recycler stash, with the Antiproton Source (Pbar) stacking, and with the MiniBooNE and Meson MCenter experiments taking beam.

Operations halted MiniBooNE beam at 9:28 AM and prepared the area for an access.

Operations prepared the Linac and Booster at 11:59 AM for safety system tests. All beam is shutdown for these tests.

At 12:08 PM, Operations reported that a Tevatron (TeV) RF station (TRF3) tripped off due to a spark. An expert investigated, but he could find nothing wrong.

At 1:52 PM, the TeV suffered a sector F4 quench.

Operations reported at 3:53 PM that safety system tests had finished and that operators had reestablished beam in Linac.

Pbar resumed stacking at 7:25 PM.

Operations established beam to MCenter and MiniBooNE at 7:33 PM.

NuMI experts began target scans at 7:41 PM.

At 9:46 PM, a Booster expert replaced a power supply (B:QFB2).

At 11:03 PM, operators turned on BRF19 per an expert's request, but it wouldn't stay on.

At 11:15 PM, a Linac RF station (KRF5) began tripping off. Operators investigated and found that LCW regulation problems had caused sparking.

At 11:22 PM, BRF2 tripped off. It reset okay.

Operators managed to get KRF5 back on and reestablished Linac beam at 11:54 PM.

At 11:59 PM, BRF8 tripped off. It reset okay.

Tuesday May 3
The midnight shift began with Operations monitoring a very small stash, with TeV experts conducting studies, with Pbar stacking, and MiniBooNE and MCenter taking beam.

At 12:35 AM, BRF8 tripped off. It reset okay.

At 12:39 AM, a Booster notcher (B:NOTCH2) tripped off. Operators called in experts who had the device reset by 12:51 AM.

TeV studies ended at 2:33 AM.

Operations put the TeV into shot setup at 3:08 AM.

At 3:22 AM, Operations reported a problem with coalesced beam in the TeV. Operators talked to an expert who suggested checking the chromaticities. Operators found that the chromaticity setting had been changed for the studies and not restored.

At 5:20 AM, BRF16 tripped off. It reset okay.

Operations established store 4125 at 5:20 AM with an initial luminosity of 72.44E30. Pbar resumed stacking.

NuMI and MCenter resumed taking beam at 5:23 AM.

At 8:58 AM, KRF5 tripped off. It reset okay.

At 9:09 AM, a Booster power supply (EAPS) tripped off when RF experts turned on BRF19. It reset okay, but then tripped off again at 9:18 AM. Experts turned it back on.

Operations reported at 10:09 AM that the EAPS trips were due to Time Line Generator (TLG) changes.

Around 10:15 AM, the Booster Notcher tripped off. Experts had the device back on ten minutes later.

At 11:19 AM, the NuMI beam permit tripped off due to a power supply limit setting (E:CPS65). Operators contacted power supply experts.

Operations put the Recycler into shot setup to stash antiprotons at 11:24 AM.

The Recycler stash ended at 12:21 PM and Operations resumed sending beam to Pbar, NuMI, and MCenter.

At 1:35 PM, a power supply expert replaced a control card for a NuMI power supply (E:CPS65). Unfortunately, this didn't help the sagging voltage problem.

At 1:55 PM, the LCW water skid for the NuMI target tripped off. It reset okay, but at 2:29 PM it tripped off again.

At 3:39 PM, the NuMI permit tripped off again, but this time due to a power loss. Experts investigated and later reported that a sump pump had failed and caused the power loss.

Operations reported at 6:23 PM that the NuMI power had been restored and that NuMI experts were beginning a study period.

The Recycler began stashing antiprotons at 9:10 PM. Operations completed the stash an hour later and then resumed sending beam to NuMI and MCenter.

Wednesday May 4
The midnight shift began with Operations monitoring store 4125 and the stash, with Pbar stacking, and with MiniBooNE, NuMI, and MCenter taking beam.

At 12:31 AM, the Booster chopper (B:CHOP) tripped off. It wouldn't reset. Operations called in experts. The experts changed a power supply and some cables.

Stacking resumed at 3:42 AM.

The Recycler began stashing antiprotons at 4:37 AM. Operations completed the stash at 5:11 AM.

Operations began preparing the Pbar rings for access at 5:26 AM.

MCenter resumed taking beam at 5:28 AM.

Operations terminated store 4125 at 6:05 AM.

Operations put the TeV into shot setup at 6:22 AM.

The Plan for Wednesday
The plan for today is to establish a new store and then to allow the safety system tests to continue throughout the day and into tomorrow.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 5/4/2005   email Fermilab